# AI on My Shoulder: Supporting Emotional Labor in Front-Office Roles with an LLM-based Empathetic Coworker

### v1

@@ 3.1. Compiling Interactions with Uncivil Clients
To design empathetic messages, we need to expose Pro-Pilot to interactions between CSRs and uncivil clients. However, actual logs of CSRs interactions tend to be protected by organizations for a variety of safety and privacy reasons. Moreover, clients from individual organizations are likely to be limited by specific scenarios. We addressed these constraints by synthetically generating a comprehensive set of client-CSR interactions. We used a real world corpus of publicly posted client complaints on Twitter (Axelbrooke, 2017), which was collected for the “study of modern customer support practices and impact.” These data served as examples to build life-like, multi-turn, text-based conversations between a CSR agent and a client seeking support over a live chat interface. The complaints in the data varied across industrial sectors, with mobiles and airlines having the largest volume.
@@ 3.1.1. Diversifying Complaints
Pro-Pilot should be robust to all kinds of complaint scenarios. To ensure our client interaction dataset captures the variety of complaints, we distilled five categories for complaints based on prior analyses of customer complaints (Cho et al., 2002)Service Quality, Product Issues, Pricing & Charges, Policy, Resolution. Four authors encoded a random sample of 15 complaints to refine the category definitions (Section A.1.1). After finalizing definitions, two authors independently encoded a random sample of 250 complaints. A third author encoded any complaints with disagreements. We also identified the sector, or domain, of each of these complaints. For the initial complaint generation, we prompt the LLM with definitions of complaint categories and a set of examples to generate a new complaint for any given input domain and category (Fig. 2).
1{Domain}: Airline
2{Category}: Product Issues
3Complaint: SouthwestAir Why would we be receiving errors when we try to check-in? Our flight takes off at 4, but we keep getting error messages.
Figure 2. The initial complaint is generated based on specified complaint ‘category‘ and organizational ‘domain‘
@@
1Phrase your responses like an UNCIVIL customer:
2- Talk in a rude, impolite, and disrespectful tone of voice.
3- Do NOT use good manners. Do NOT use courtesy.
4- Act with disregard to others.
Figure 3. An abridged version of the prompt for the uncivil Client-Agent
@@
Prompt given to Pro-Pilot to simulate an uncivil interaction.
@@ 3.1.2. Prompting for Incivility
Once a complaint was initiated, we created a distinct Client-Agent to carry forward a turn-by-turn interaction with a CSR. Recent explorations with LLMs show that chat agents can be infused with specific personalities and reliably respond to fit the specified personality (Serapio-García et al., 2023). Uncivil clients are likely to induce stressful scenarios in emotional labor (Andersson and Pearson, 1999). We refer to Andersson and Pearson’s definition of incivility to design a role for Client-Agent (Fig. 3). To respond to the client, we set up another agent, Rep-Agent, to act as “a service representative chatting with a customer online.” Both agents included an intermediate step to contextualize the input with respect to the conversation history so that the subsequent response maintains continuity. By having Client-Agent and the Rep-Agent converse, we were able to design multi-turn conversations where Client-Agent is acting uncivil to the Rep-Agent. We refer to these synthetic, multi-turn conversations as incidents. To capture the variety of scenarios in front-office work, we varied incidents across 3 domains — airlines and hotels, which have been used as representative scenarios in the literature (Frey-Cordes et al., 2020), and mobiles, which are extensively represented in the real world complaint dataset (Axelbrooke, 2017). For each domain and each complaint category, we generated 3 incidents. Each incident contained 5 total turns, with the Client-Agent awaiting a response to their last message. The incomplete conversation was purposefully intended to assess how Pro-Pilot could digest ongoing incidents and provide support.
@@ 3.2. Emotional Regulation with Empathy
A coworker with high emotional intelligence can mitigate the client incivility by helping CSRs regulate their emotions (Yang and Lau, 2019). The primary component of Pro-Pilot, Emo-Reframe, was designed to reflect this social phenomenon.
@@ 3.2.1. Examples of Human Reframing
Prior work has shown that LLMs can help individuals perform self-directed ER (Sharma et al., 2023). We leveraged Sharma et al.’s dataset  [^2] This dataset was developed with mental health practitioners on cognitive restructuring and emotional reframing to seed examples of our own implementation (Sharma et al., 2023). This dataset contains tuples of a situation that triggers a negative thought and a corresponding reframed thought for ER. Client incivility differentiates itself from everyday negative thinking, which often stems from abstract concerns. By contrast, incivility is interpersonal in nature and causes an ego threat—an attack on one’s self-esteem and image leading to retaliatory thoughts (Frey-Cordes et al., 2020). Thus, we identified situations in the dataset that were interpersonal, involved at least two individuals (e.g., “I was talking to a friend who got me angry”) and described a confrontation, in-the-moment conflict (e.g., “I get so annoyed and frustrated when my baby cries”). Moreover, we narrowed down examples that labeled retaliatory negative thoughts, such as blaming and labeling  [^3] These categories were already annotated in the original dataset, and we defined an inclusion criterion along these.
@@ 3.2.2. Reframing Chain-of-Thought
Since the dataset was designed around self-directed ER, the outputted reframe message goes through a rudimentary prompt to rephrase it as a message addressed to the CSR.
4. Technical Evaluation: Inspecting the Efficacy and Perception of Pro-Pilot
Our central metaphor for the design of Pro-Pilot is the support of human coworkers (Yang and Lau, 2019). Therefore, before embedding Pro-Pilot into a real-time interaction, we first evaluated the quality of the support messages that Emo-Reframe produces. To answer RQ I, we analyzed the messages to provide empirical, statistical evidence on the differences and similarities between Pro-Pilot and a human coworker.
Table 1. Sample client-incivility incident from our study (domain: Mobiles and category: Product Issues)
@@ 4. Technical Evaluation: Inspecting the Efficacy and Perception of Pro-Pilot
Pro-Pilot Hey, remember that the customer’s frustration is about their situation, not you. Stay calm, ask clear questions to get all the info you need, and reassure them you’re dedicated to fixing the issue swiftly. Keep your professionalism and empathy at the forefront to manage the situation effectively and offer the best help.
@@
4.1. Method: Comparing Pro-Pilot to Coworker Empathy
@@
Phase I - Writing Empathetic Messages for CSRs: Participants read incidents of client incivility (Section 3.1.2) and provided an empathetic message as if they were a coworker. Eligible participants needed to have at least 1 year of relevant experience, have encountered uncivil clients, and primarily interact with clients via computer-mediated communication (e.g., live chat). Each participant was randomly assigned a scenario (airlines, hotel, or mobile devices) and they viewed 6 incidents (at least one of each complaint category). Participants were guided to write an empathetic message using Downward-Arrow-Technique (DAT) (Burger et al., 2021) that leads them to first consider the CSR’s emotion, then their thought, and finally describe a message to overcome the thought. Their messages are indicated by the notation, Human.
@@
To foster more variety in the task, we injected contextual information about the CSR in some of the incidents. Research on workplaces has revealed behavioral indicators of worker mental state, such as focused, stressed, and bored (Mark et al., 2014; Morshed et al., 2022). Based on these studies we randomly assigned descriptions of CSR mental state for 2 incidents (Section A.1.2). As per organizational psychology, personality can serve as an important context to explain how workers perceive situations (Donnellan and Robins, 2010). Participants were asked to think of an actual coworker and select their personality as either resilient (organized and dependable), undercontrolled (competitive and high energy), or overcontrolled (detail-oriented) archetype (Donnellan and Robins, 2010). 2 incidents included descriptions of these personalities (Section A.1.3). 116 CSRs successfully completed the writing task and they received \$$\$8$ for the $30$ minutes of their time.
@@
Phase II - Scoring Empathetic Messages from coworkers: Participants read incidents of client incivility and rated the perceived empathy of multiple empathetic support messages. They were not explicitly informed which messages originated from Pro-Pilot and which ones were from humans. To represent multiple dimensions of empathy, we included five dimensions drawn from the literature — sincerity, compassion, warmth, actionability, and relatability (Sharma et al., 2023; Nambisan, 2011). Participants were screened using the same criteria but did not overlap with the previous task. Participants rated these on a 7-point semantic scale. 143 CSRs evaluated the 10 incidents each and were compensated \$$\$5$ for the $15$ minute task.
@@
Analysis: The inherent characteristics of language are associated with its effectiveness in communication. Research on linguistics has revealed certain linguistic attributes that are important to explain social support and psychotherapy (Norcross and Lambert, 2018). More recent works have also stressed the importance of language in human-AI interaction (Wang et al., 2021). In fact, contemporary studies have already validated linguistic models of empathy (Buechel et al., 2018; Tafreshi et al., 2021). Our analyses include such models and go beyond to measure additional metrics that help interpret Pro-Pilot’s messages in light of Human’s messages (retrieved from Phase I). First, we tested how easily CSRs can read and comprehend messages (Syntax and Structure), because these attributes can determine how meaningful the support message is to the reader (Glass and Maddox, 1992; Mcinnes and Haglund, 2011; Blum-Kulka and Olshtain, 1986). Second, we measured the style and meaning of messages to capture social aspects of communication (Linguistic Style & Semantics). The attributes here capture more colloquial conceptions of human empathy (Pennebaker et al., 2014; Zhou et al., 2021; Sharma et al., 2023). Further, we inspected the words that were used in the messages to identify relevant psycholinguistic markers that reflect social support (Tausczik and Pennebaker, 2010). The differences we measured provide a linguistic landscape of messages, but empathy is highly contextual and its appropriateness varies by situation (Zhou et al., 2021). Consequently, we also compared the differences in perceived empathy scores from ratings in Phase II. For every comparison we performed a paired $t$-test and computed the effect sizes of differences using Cohen’s $d$.
4.2. Lexico-Semantic Analyses
@@
We operationalized the different lexico-semantic aspects of messages based on domain literature to understand the differences between Pro-Pilot and Human. Table 3 describes all the key results. We only report on statistically significant and theoretically relevant metrics. Below, we describe the theory-driven rationale behind the choice and operationalization of the measures, along with our observations.
@@
Verbosity and Repeatability  The length and thoroughness of messages explain their effectiveness in providing support (Saha and Sharma, 2020; Glass and Maddox, 1992). The richness of expressions in communication can be described using verbosity and repeatability (Kolden et al., 2011). Verbosity describes the level of detail and conciseness in supportive communication. We operationalized verbosity as the number of words per thought reframing. Repeatability accounts for the reuse of words in a piece of text. Higher repeatability indicates a lack of conciseness. Drawing on prior work (Ernala et al., 2017; Saha and Sharma, 2020; Yuan et al., 2023), we operationalized repeatability as the normalized occurrence of non-unique words.  Table 3 shows the statistically significant differences in both verbosity and repeatability. Pro-Pilot’s reframings were 68% longer (Cohen’s $d$=1.60) and 36% more repeatable (Cohen’s $d$=1.54) than Human messages. Sociolinguistic theory argues that the use of more words can indicate sincerity and effort in putting one’s point across (Blum-Kulka and Olshtain, 1986). Having said that, Pro-Pilot’s verbose messages might not always be compatible with the urgency of certain demanding conversations.
Readability.  Apart from the shape of the message, the vocabulary and style can also determine how easily it can be read. Therefore, we turn to the measure of readability, to understand how Pro-Pilot’s messages might be comprehended (Wang et al., 2013; Mcinnes and Haglund, 2011). According to Wang et al. people perceive AI as more intelligent when the readability is higher (Wang et al., 2021). Per prior work (Wang et al., 2021; Saha and Sharma, 2020; Pitler and Nenkova, 2008), we utilized the Coleman-Liau Index ($\mathtt{CLI}$(Coleman and Liau, 1975), which assesses readability based on a sentence’s character and word structure. $\mathtt{CLI}$ approximates the U.S. grade-level required to read certain text. It is operationalized as follows: $\mathtt{CLI=(0.0588L-0.296S-15.8)}$, where $\mathtt{L}$ is the average number of letters per 100 words, and $\mathtt{S}$ is the average number of sentences per 100 words.
The readability of Pro-Pilot’s thought reframings was 36% higher than that of Human’s reframings (Cohen’s $d$=1.54). Although a higher $\mathtt{CLI}$ score by Pro-Pilot indicates better writing quality, it also implies a more advanced level of English education may be required to fully understand its content (Coleman and Liau, 1975). While Human content only needs an average of 12.05 years of education, Pro-Pilot’s content required about 16.44 years. Thus, Pro-Pilot might be more effective for some CSRs but difficult to comprehend for others.

### v2

@@ 3.1. Compiling Interactions with Uncivil Clients
To design empathetic messages, we need to expose Care-Pilot to interactions between CSRs and uncivil clients. However, actual logs of CSRs interactions tend to be protected by organizations for a variety of safety and privacy reasons. Moreover, clients from individual organizations are likely to be limited by specific scenarios. We addressed these constraints by synthetically generating a comprehensive set of client-CSR interactions. We used a real world corpus of publicly posted client complaints on X (formerly known as Twitter) (Axelbrooke, 2017), which was collected for the “study of modern customer support practices and impact.” These data served as examples to build life-like, multi-turn, text-based conversations between a CSR agent and a client seeking support over a live chat interface. The complaints in the data varied across industrial sectors, with mobiles and airlines having the largest volume. We provide the complete prompts for this element of our system in Appendix A.
@@ 3.1.1. Diversifying Complaints
Care-Pilot should be robust to all kinds of complaint scenarios. To ensure our client interaction dataset captures the variety of complaints, we distilled five categories for complaints based on prior analyses of customer complaints (Cho et al., 2002)Service Quality, Product Issues, Pricing & Charges, Policy, Resolution. Four authors encoded a random sample of 15 complaints to refine the category definitions (Appendix A.1). After finalizing definitions, two authors independently encoded a random sample of 250 complaints. A third author encoded any complaints with disagreements. We also identified the sector, or domain, of each of these complaints. For the initial complaint generation, we prompt the LLM with definitions of complaint categories and a set of examples to generate a new complaint for any given input domain and category. (Fig. 2) shows one such example. Fig. A3 shows a set of examples we randomly sampled from our identified examples to ensure variety of category and domain.
Figure 2. The Client-Agent learns different types of complaints through examples. The initial complaint is generated based on specified complaint ‘category‘ and organizational ‘domain‘
@@
Figure 3. Abridged version of the prompt for the uncivil Client-Agent describes the details of the role needed to simulate an challenging client.
@@
Prompt given to Care-Pilot to simulate an uncivil interaction.
@@ 3.1.2. Prompting for Incivility
Once a complaint was initiated, we created a distinct Client-Agent to carry forward a turn-by-turn interaction with a CSR. Recent explorations with LLMs show that chat agents can be infused with specific personalities and can reliably respond to fit the specified personality (Serapio-García et al., 2023). Uncivil clients are likely to induce stressful scenarios in emotional labor (Andersson and Pearson, 1999). We refer to Andersson and Pearson’s definition of incivility to design a role for Client-Agent (Fig. 3). In the appendix, Fig. A1 details the complete prompt for uncivil behavior that was crafted to reduce hallucinations (“Do NOT reveal your role”) and ensure closure of conversation (“After 12 turns , do NOT respond further..”). To respond to the client, we set up another agent, Rep-Agent, to act as “a service representative chatting with a customer online.” Both agents included an intermediate step to contextualize the input with respect to the conversation history so that the subsequent response maintains continuity (Fig. A2). By having Client-Agent and the Rep-Agent converse, we were able to design multi-turn conversations where Client-Agent is acting uncivil to the Rep-Agent. We refer to these synthetic, multi-turn conversations as incidents. To capture the variety of scenarios in front-office work, we varied incidents across three domains — airlines and hotels, which have been used as representative scenarios in the literature (Frey-Cordes et al., 2020), and mobiles, which are extensively represented in the real world complaint dataset (Axelbrooke, 2017). For each domain and each complaint category, we generated 3 incidents, resulting in a total of $45$ unique incidents. These incidents can be retrieved from the supplementary data. Each incident contained 5 total turns, with the Client-Agent awaiting a response to their last message. The incomplete conversation was purposefully intended to assess how Care-Pilot could digest ongoing incidents and provide support. These synthetically generated client incivility incidents are also in our supplementary data.
@@ 3.2. Emotional Regulation with Empathy
A coworker with high emotional intelligence can mitigate the client incivility by helping CSRs regulate their emotions (Yang and Lau, 2019). The primary component of Care-Pilot, Emo-Reframe, was designed to reflect this social phenomenon.
@@ 3.2.1. Examples of Human Reframing
Prior work shows that LLMs can help individuals perform self-directed ER (Sharma et al., 2023). We leveraged Sharma et al.’s dataset  [^3] This dataset was developed with mental health practitioners on cognitive restructuring and emotional reframing to seed examples of our own implementation with few-shot learning (Wang et al., 2020; Sharma et al., 2023). This dataset contains tuples of a situation that triggers a negative thought and a corresponding reframed thought for ER. Client incivility differentiates itself from everyday negative thinking, which often stems from abstract concerns. By contrast, incivility is interpersonal in nature and causes an ego threat—an attack on one’s self-esteem and image leading to retaliatory thoughts (Frey-Cordes et al., 2020). Thus, we identified situations in the dataset that were interpersonal, involved at least two individuals (e.g., “I was talking to a friend who got me angry”) and described a confrontation, in-the-moment conflict (e.g., “I get so annoyed and frustrated when my baby cries”). Moreover, we narrowed down examples that labeled retaliatory negative thoughts, such as blaming and labeling  [^4] These categories were already annotated in the original dataset, and we defined an inclusion criterion along these. Our selected examples are included in the supplementary data.
@@ 3.2.2. Reframing Chain-of-Thought
Since the dataset was designed around self-directed ER, the outputted reframe message goes through a rudimentary prompt to rephrase it as a message addressed to the CSR. The initial portion of Fig. 1(a) illustrates how reframing occurs for a given incident. Moreover, in Appendix B describes the exact prompt we used for this study and the supplementary data includes all possible reframing examples that we deemed appropriate for Care-Pilot.
4. Technical Evaluation: Inspecting the Efficacy and Perception of Care-Pilot
Our central metaphor for Care-Pilot is the support of human coworkers (Yang and Lau, 2019). Therefore, before embedding Care-Pilot into a real-time interaction, we first evaluated the quality of messages that Emo-Reframe produces. To answer RQ I, we analyzed the messages to provide empirical, statistical evidence on the differences and similarities between Care-Pilot and a human coworker.
Table 1. Sample client-incivility incident from our study (domain: Airlines and category: Product Issues)
@@ 4. Technical Evaluation: Inspecting the Efficacy and Perception of Care-Pilot
Care-Pilot Hey, remember that the customer’s frustration is about their situation, not you. Stay calm, ask clear questions to get all the info you need, and reassure them you’re dedicated to fixing the issue swiftly. Keep your professionalism and empathy at the forefront to manage the situation effectively and offer the best help.
@@
4.1. Method: Comparing Care-Pilot to Coworker Empathy
@@
Phase I — Writing Empathetic Messages for CSRs: Participants read incidents of client incivility (Section 3.1.2) and provided an empathetic message as if they were a coworker. Eligible participants needed to have at least 1 year of relevant experience, have encountered uncivil clients, and primarily interact with clients via computer-mediated communication (e.g., live chat). Each participant was randomly assigned a scenario (airlines, hotel, or mobile devices) and they viewed 6 incidents (at least one of each complaint category). Participants were guided to write an empathetic message using Downward-Arrow-Technique (DAT) (Burger et al., 2021). This technique is used in cognitive psychotherapy to help people reflect on emotions by describing a causal schema of maladaptive emotions by successive instrospective questions. Prior studies involving datasets of emotion labeling of situations have used this method (Burger et al., 2021; Sharma et al., 2023). Our adaption of DAT led participants to first describe the CSR’s emotion with 1 word; then identify the closest emotion among (sadness, fear, anger, happiness); next describe the CSR’s thought behind the emotion (1-2 lines); finally describe a message to overcome the thought. Their messages are indicated by the notation, Human.
@@
To foster more variety in the task, we injected contextual information about the CSR in some incidents. Research on workplaces has revealed behavioral indicators of worker mental state, such as focused, stressed, and bored (Mark et al., 2014; Morshed et al., 2022). Based on these studies, we randomly assigned descriptions of CSR mental state for 2 incidents (Appendix C.1). As per organizational psychology, personality can serve as an important context to explain how workers perceive situations (Donnellan and Robins, 2010). Participants were asked to think of an actual coworker and select their personality as either resilient (organized and dependable), undercontrolled (competitive and high energy), or overcontrolled (detail-oriented) archetype (Donnellan and Robins, 2010). 2 incidents included descriptions of these personalities (Appendix C.2). 116 CSRs successfully completed the writing task, and they received \$$\$8$ for the $30$ minutes of their time.
@@
Phase II — Scoring Empathetic Messages from coworkers: Participants read incidents of client incivility and rated the perceived empathy of multiple empathetic support messages. To represent multiple aspects of empathy, we included five dimensions drawn from the literature (Sharma et al., 2023; Nambisan, 2011):
  • Sincerity: The genuineness or authenticity of the expression of concern.
  • Compassion: The awareness of others’ challenges and a comprehension of where their difficulties arise from.
  • Warmth: The approachability or comfort exhibited by the tone of the message.
  • Actionability: The offering of practical assistance.
  • Relatability: The extent to which the message aligns with the context of another.
Participants were screened using the same criteria but did not overlap with the previous task. They were asked to review client–representative conversations and evaluate coworkers’ empathetic messages. The origin of the message, whether Care-Pilot or Human, was not explicitly revealed. Table 2 shows different messages from different sources, but referring to the same incident. After viewing the incident, participants were first asked to select an emotion the CSR would be feeling (e.g., ashamed, attentive, bullied, curious, disconnected, resolute, and rushed   [^5] The list of emotions participants could select from were the unique emotions reported by participants in Phase I using the DAT method. We list these emotions in Appendix C.3). Then they evaluated the different messages with the prompt, “Evaluate the effectiveness of the message below in helping the representative overcome their feeling.” Each message was followed by five 7-point semantic differential scales representing each of the above dimensions (e.g., insincere/sincere and cold/warm). 143 CSRs were tasked to evaluate 6 incidents each and were compensated \$$\$5$ for the $15$ minute task.
@@
Analysis: The inherent characteristics of language are associated with its effectiveness in communication. Research on linguistics has revealed certain linguistic attributes that are important to explain social support and psychotherapy (Norcross and Lambert, 2018). More recent works have also stressed the importance of language in human-AI interaction (Wang et al., 2021). In fact, contemporary studies have already validated linguistic models of empathy (Buechel et al., 2018; Tafreshi et al., 2021). Our analyses include such models and go beyond to measure additional metrics that help interpret Care-Pilot’s messages in light of Human’s messages (retrieved from Phase I). While we had $45$ unique client-complaint incidents (Section 3.1.2), we also had additonal contextual variations for personality and behavior. Therefore, our sample included $315$ unique variations. Collectively, our participants produced $660$ different empathetic messages corresponding to different incident variations. These messages (Human) were paired with Care-Pilot’s empathetic messages for the same incident variations. Following this, we computed a variety of domain-driven linguistic attributes. First, we tested how easily CSRs can read and comprehend messages (Syntax and Structure), because these attributes can determine how meaningful the support message is to the reader (Glass and Maddox, 1992; Mcinnes and Haglund, 2011; Blum-Kulka and Olshtain, 1986). Second, we measured the style and meaning of messages to capture social aspects of communication (Linguistic Style & Semantics). The attributes here capture more colloquial conceptions of human empathy (Pennebaker et al., 2014; Zhou et al., 2021; Sharma et al., 2023). Further, we inspected the words that were used in the messages to identify relevant psycholinguistic markers that reflect social support (Tausczik and Pennebaker, 2010). The differences we measured provide a linguistic landscape of messages, but empathy is highly contextual and its appropriateness varies by situation (Zhou et al., 2021). Consequently, we also compared the differences in perceived empathy scores from ratings in Phase II. The messages in this phase were paired up as they were shown to participants and we maintained the same pairs while statistically comparing the differences. Some participants only completed the task partially, which amounted to $552$ pairs for us to compare. The supplementary files include the data collected from both phases along with our computed and retrieved scores for the messages. For every comparison we performed a paired $t$-test and computed the effect sizes of differences using Cohen’s $d$. Our supplementary data includes the Human and Care-Pilot messages along with their measured and human-annotated metrics that we used for analyses.
4.2. Findings: Lexico-Semantic Analyses
@@
We operationalized the different lexico-semantic aspects of messages based on domain literature to understand the differences between Care-Pilot and Human. Table 3 describes all the key results. We only report on statistically significant and theoretically relevant metrics. Below, we describe the theory-driven rationale behind the choice and operationalization of the measures, along with our observations.
@@
Categories $\mu$ (CPilot) $\mu$ (Hum.) Diff % Cohen’s $d$ $t$-test
Lexico-Semantics
Syntax & Structure
Verbosity 57.46 34.29 67.56 1.60   28.72 ***
Repeatability 0.20 0.13 55.32 1.17   21.53 ***
Readability 16.44 12.05 36.40 1.54   28.03 ***
Style & Semantics
Categorical Dynamic Index (CDI) 14.81 3.89 281.14 0.87   16.10 ***
Empathy 0.91 0.90 1.85 0.40   7.15 ***
Emotional Reactivity 1.00 0.89 12.52 0.31   0.19 ***
Adaptability 0.81 0.77 4.96 0.67   15.28 ***
Psycholinguistics
Affect
Pos. Affect 0.043 0.049 -11.55   -0.16 -2.93 **
Anger 0.020 0.012 61.77 0.47   8.51 ***
Sad 0.001 0.004 -74.87   -0.31 -5.62 ***
Interpersonal Focus (Pronouns)
1st P. Sin. 0.005 0.015 -67.36   -0.58 -10.52 ***
1st P. Plu. 0.002 0.007 -75.43   -0.37 -6.77 ***
2nd P. 0.051 0.062 -17.20   -0.32 -5.79 ***
@@
Verbosity and Repeatability  The length and thoroughness of messages explain their effectiveness in providing support (Saha and Sharma, 2020; Glass and Maddox, 1992). The richness of expressions in communication can be described using verbosity and repeatability (Kolden et al., 2011). Verbosity describes the level of detail and conciseness in supportive communication. We operationalized verbosity as the number of words per thought reframing. Repeatability accounts for the reuse of words in a piece of text. Higher repeatability indicates a lack of conciseness. Drawing on prior work (Ernala et al., 2017; Saha and Sharma, 2020; Yuan et al., 2023), we operationalized repeatability as the normalized occurrence of non-unique words. Table 3 shows the statistically significant differences in both verbosity and repeatability. Pro-Pilot’s reframing messages were 68% longer (Cohen’s $d$=1.60) and 36% more repeatable (Cohen’s $d$=1.54) than Human messages. Sociolinguistic theory argues that the use of more words can indicate sincerity and effort in putting one’s point across (Blum-Kulka and Olshtain, 1986). Having said that, Care-Pilot’s verbose messages might not always be compatible with the urgency of certain demanding conversations.
Readability.  Apart from the shape of the message, vocabulary and style also determine the ease of reading. Therefore, we turn to the measure of readability, to understand how Care-Pilot’s messages might be comprehended (Wang et al., 2013; Mcinnes and Haglund, 2011). According to Wang et al. people perceive AI as more intelligent when the readability is higher (Wang et al., 2021). Per prior work (Wang et al., 2021; Saha and Sharma, 2020; Pitler and Nenkova, 2008), we utilized the Coleman-Liau Index ($\mathtt{CLI}$(Coleman and Liau, 1975), which assesses readability based on a sentence’s character and word structure. $\mathtt{CLI}$ approximates the U.S. grade-level required to read certain text. It is operationalized as follows: $\mathtt{CLI=(0.0588L-0.296S-15.8)}$, where $\mathtt{L}$ is the average number of letters per 100 words, and $\mathtt{S}$ is the average number of sentences per 100 words.
The readability of Pro-Pilot’s thought reframing messages were 36% higher than that of Human’s reframing messages (Cohen’s $d$=1.54). Although a higher $\mathtt{CLI}$ score by Pro-Pilot indicates better writing quality, it also implies a more advanced level of English education may be required to fully understand its content (Coleman and Liau, 1975). While Human content only needs an average of 12.05 years of education, Pro-Pilot’s content required about 16.44 years. Thus, Care-Pilot might be more effective for some CSRs but difficult to comprehend for others.